OpenStudy (anonymous):

If f(x) = 12x + 7, what is f^-1(x)

OpenStudy (ash2326):

One important thing to note \[f^{-1} (f (x))=f(f^{-1} (x))=x\] here we are given f(x) \[f(x)=12x+7\] Put \(x=f^{-1} (x)\) We get \[f(f^{-1} (x))= 12 \times f^{-1} (x)+7\] so we get \[x=12 \times f^{-1} (x)+7\] Could you find \(f^{-1} (x)\) now?

OpenStudy (anonymous):

In other words, y=12x+7. Switch x and y and solve for y. x=12y+7.